# A Differentiated Reward Method for Reinforcement Learning based Multi-Vehicle Cooperative Decision-Making Algorithms

### v1

@@ V-B Compared Methods
We compare the training and deployment performance of widely used multi-agent reinforcement learning algorithms, including MADQN [23], MAPPO [27], and QMIX [28], by employing the generally adopted vehicle decision reward function (Equation (10)), the centering reward function (oracle centering in [19]), and the differentiated reward function (Equation (14)).
@@
Learning Rate $3\times 10^{-4}$
@@
Target Network Update Interval 100 steps
Exploration Rate Decay 0.99 (linear)
@@ V-C Evaluation Metric
Avg. Speed, reflecting overall traffic efficiency:
$\bar{v}=\frac{\sum_{t=0}^{\text{END}}\bar{v}_{t}}{\sum_{t=0}^{\text{END}}1},$
@@
Min. Gap, quantifying potential collision risk:
$C_{\text{GAP}}=\frac{\sum_{e=0}^{N_{\text{epi}}}g_{\text{min, }e}}{N_{\text{epi}}},$
@@
Lane Change Frequency, evaluating the rationality of lane resource utilization:
$F_{\text{LC}}=\frac{\sum_{t=0}^{\text{END}}n_{\text{LC},t}}{\sum_{t=0}^{\text{END}}n_{\text{CAV},t}},$
@@
Succ. Rate:
@@
TABLE III: Metrics comparison of reward functions in Qmix
Pene. Rate Rew. Fn. Avg. Speed (m/s) Min. Gap(m) LC. Rate (time/min) Succ. Rate
25% GR 7.95±0.4 1.33±0.8 0.67±0.03 49.20±0.2
CR 10.95±0.3 5.36±0.6 0.14±0.02 61.45±0.1
DR 11.46±0.2 36.47±0.4 0.07±0.02 96.49±0.1
50% GR 1.78±0.5 4.91±1.2 0.01±0.05 36.21±0.3
CR 14.18±0.4 5.55±0.9 0.10±0.04 68.83±0.2
DR 12.06±0.3 49.04±0.7 0.04±0.03 97.01±0.1
70% GR 0.16±0.5 21.88±1.2 0.04±0.05 40.09±0.3
CR 0.55±0.4 6.01±0.9 0.07±0.04 70.25±0.2
DR 12.01±0.3 53.56±0.7 0.01±0.03 94.64±0.1
100% GR 0.84±0.5 6.07±1.2 0.01±0.05 30.29±0.3
CR 12.58±0.4 5.24±0.9 0.15±0.04 62.88±0.2
DR 16.79±0.3 32.30±0.7 0.10±0.03 88.82±0.1
@@ V-D Results and Comparison
Figure 3 illustrates the training process of the algorithms. To better capture the early-stage convergence and later-stage stability, a logarithmic scale was applied to the horizontal axis.
@@
It can be observed that in MADQN, the training curves for all three reward functions exhibit significant fluctuations, indicating poor stability. Despite tuning the algorithm to its optimal state, the performance of the three reward functions varies across different penetration rates. This instability is likely due to MADQN being based on independent Q-learning, which struggles to adapt to multi-vehicle cooperative decision-making tasks. Similarly, in MAPPO, although the algorithm achieves stable convergence, the performance of the centralized reward function exhibits considerable deviation in the stabilized phase under the same settings. Specifically, for penetration rates of 0.25, 0.5, and 1.0, the orange curve shows substantial deviation after 20,000 episodes. As a result, in Table III, we only provide quantitative performance metric analyses for QMIX.
@@
In the experiments with MAPPO and QMIX, it is evident that the convergence speed of the differential reward function is significantly faster than that of the centralized and general reward functions. Additionally, the return values of the centralized reward function are noticeably higher than those of the general reward function, particularly in the later training stages (>20,000 episodes). This finding is consistent with the conclusions reported in [19].
@@
Table III shows the performance metrics of the 3 reward functions under the QMIX algorithm at different penetration rates. It can be observed that, in most cases, the differentiated reward function achieves the best performance in terms of traffic efficiency, safety, action rationality, and task completion rate.
@@ VI Conclusion and future work
This paper proposes a differentiated reward method for RL based multi-vehicle cooperative decision-making algorithms. By incorporating state transition gradient information into the reward function design, the method resolves the issue of distinguishing action values in steady-state traffic flow, a challenge commonly encountered with traditional reward mechanisms. Experimental results shows that the differentiated reward method significantly improves the training efficiency and decision quality of multi-agent algorithms, outperforming centering and general reward functions in core metrics such as traffic efficiency, safety, and action rationality. Moreover, the method maintains stable performance across different autonomous driving penetration rates (25%-100%), which shows good scalability and learning stability of our method. Future research directions include further optimization of reward function design, exploration of more complex traffic scenarios, and large-scale validation and application in real-world traffic systems.

### v2

@@ V-A Simulation environment and experiment settings
Maximum Speed 25 m/s
$d_{\text{max}}$ in Equation (1)
Longitudinal Observation Range $\pm$ 100m
Lateral Observation Range $\pm$ 1 adjacent lanes
@@ V-B Compared Methods
We compare the training and deployment performance of widely used multi-agent reinforcement learning algorithms, including MADQN [23], MAPPO [27], and QMIX [28], by employing the generally adopted vehicle decision reward function (GR, with Equation (18)), the centering reward function (CR, with oracle centering in [19]), and the differentiated reward function (DR, with Equation (23)).
@@
Total Episodes 50000
Episode Duration 18 s (180 steps)
RMSProp Learning Rate $3\times 10^{-4}$
RMSProp $\alpha$ 0.99
RMSProp $\epsilon$ $1\times 10^{-5}$
Batch Size 32 (episodic)
@@
Target Network Update Interval 10 episodes
Exploration Rate Decay 0.998 (linear, episodic)
Weights $\omega_{1,2,3,4}$ in Equation (19) 10, 1e3, 1, -5
Reward for Arrival 30
$\sigma$, $\zeta$ in Equation (23)
TABLE III: Metrics comparison of reward function in QMIX
Pene. Rate Rew. Fn. Avg. Speed (m/s) Min. Gap (m) LC. Freq. (time/min) Succ. Rate (%)
25% GR 7.95$\pm$0.42 1.33$\pm$0.74 0.67$\pm$0.03 49.20$\pm$0.24
CR 10.95$\pm$0.33 5.36$\pm$0.62 0.14$\pm$0.02 61.45$\pm$0.15
DR 11.46$\pm$0.18 36.47$\pm$0.36 0.07$\pm$0.02 96.49$\pm$0.12
50% GR 7.19$\pm$0.83 4.26$\pm$0.55 0.14$\pm$0.01 33.72$\pm$2.66
CR 14.18$\pm$0.41 5.55$\pm$0.91 0.10$\pm$0.04 68.83$\pm$0.19
DR 12.06$\pm$0.28 49.04$\pm$0.68 0.04$\pm$0.03 97.01$\pm$0.14
75% GR 6.31$\pm$1.37 6.01$\pm$4.51 0.12$\pm$0.04 42.16$\pm$5.60
CR 10.60$\pm$2.44 5.22$\pm$1.87 0.11$\pm$0.45 73.78$\pm$9.18
DR 12.01$\pm$0.35 53.56$\pm$0.74 0.01$\pm$0.00 94.64$\pm$0.08
100% GR 11.89$\pm$1.07 6.76$\pm$1.87 0.21$\pm$0.08 34.37$\pm$2.74
CR 12.58$\pm$0.42 5.24$\pm$0.86 0.15$\pm$0.04 62.88$\pm$0.23
DR 16.79$\pm$0.33 32.30$\pm$0.72 0.10$\pm$0.03 88.82$\pm$0.07
[…]
@@ V-C Evaluation Metric
Average Speed (Avg. Speed), reflecting overall traffic efficiency:
$\bar{v}=\frac{\sum_{t=0}^{\text{END}}\bar{v}_{t}}{\sum_{t=0}^{\text{END}}1},$ (26)
@@
Minimum Headway (Min. Gap), quantifying potential collision risk:
$C_{\text{GAP}}=\frac{\sum_{e=0}^{N_{\text{epi}}}g_{\text{min, }e}}{N_{\text{epi}}},$ (27)
@@
Lane Change Frequency (LC. Freq.), evaluating the rationality of lane resource utilization:
$F_{\text{LC}}=\frac{\sum_{t=0}^{\text{END}}n_{\text{LC},t}}{\sum_{t=0}^{\text{END}}n_{\text{CAV},t}},$ (28)
@@
Mission Success Rate (Succ. Rate):
@@
TABLE IV: Metrics comparison of reward function in MADQN
Pene. Rate Rew. Fn. Avg. Speed (m/s) Min. Gap (m) LC. Freq. (time/min) Succ. Rate (%)
25% GR 4.38$\pm$1.20 1.85$\pm$1.64 0.34$\pm$0.06 30.15$\pm$2.49
CR 4.88$\pm$1.65 4.05$\pm$2.23 0.23$\pm$0.07 36.61$\pm$3.45
DR 4.44$\pm$1.03 17.00$\pm$2.68 0.37$\pm$0.02 56.00$\pm$2.14
50% GR 3.09$\pm$1.11 2.94$\pm$1.24 0.32$\pm$0.07 24.08$\pm$1.29
CR 5.57$\pm$2.64 6.50$\pm$2.37 0.34$\pm$0.10 38.14$\pm$3.63
DR 4.66$\pm$2.13 28.99$\pm$0.75 0.28$\pm$0.03 43.75$\pm$1.16
75% GR 5.67$\pm$3.33 6.42$\pm$3.79 0.38$\pm$0.12 35.86$\pm$1.04
CR 6.32$\pm$1.80 6.12$\pm$2.29 0.40$\pm$0.09 40.96$\pm$2.32
DR 13.44$\pm$3.16 17.16$\pm$0.75 0.36$\pm$0.13 47.73$\pm$2.18
100% GR 4.70$\pm$2.53 4.38$\pm$2.04 0.16$\pm$0.06 21.66$\pm$3.47
CR 7.61$\pm$3.50 7.10$\pm$4.64 0.45$\pm$0.15 46.07$\pm$2.87
DR 9.89$\pm$2.48 15.10$\pm$0.72 0.34$\pm$0.03 39.96$\pm$0.18
TABLE V: Metrics comparison of reward function in MAPPO
Pene. Rate Rew. Fn. Avg. Speed (m/s) Min. Gap (m) LC. Freq. (time/min) Succ. Rate (%)
25% GR 4.05$\pm$0.58 8.94$\pm$0.69 0.42$\pm$0.07 30.97$\pm$4.55
CR 5.54$\pm$1.45 16.63$\pm$0.14 0.29$\pm$0.03 28.32$\pm$4.13
DR 8.18$\pm$3.19 14.56$\pm$0.49 0.03$\pm$0.02 72.43$\pm$0.15
50% GR 4.06$\pm$1.71 10.35$\pm$0.36 0.48$\pm$0.03 35.94$\pm$3.39
CR 5.63$\pm$3.99 16.28$\pm$0.16 0.05$\pm$0.01 41.50$\pm$1.36
DR 13.36$\pm$2.74 41.18$\pm$0.76 0.08$\pm$0.03 56.03$\pm$0.13
75% GR 4.44$\pm$2.43 12.71$\pm$0.19 0.21$\pm$0.06 23.42$\pm$3.67
CR 7.57$\pm$3.27 18.20$\pm$0.43 0.14$\pm$0.02 41.11$\pm$1.25
DR 14.27$\pm$2.36 49.51$\pm$0.75 0.11$\pm$0.03 60.06$\pm$0.16
100% GR 4.80$\pm$2.82 13.11$\pm$0.22 0.37$\pm$0.14 33.46$\pm$4.95
CR 6.43$\pm$4.03 27.65$\pm$0.17 0.12$\pm$0.02 29.69$\pm$1.24
DR 18.47$\pm$2.82 46.15$\pm$0.72 0.04$\pm$0.03 66.93$\pm$0.17
@@ V-D Results and Comparison
Fig. 3 illustrates the training process of the algorithms. To better capture the early-stage convergence and later-stage stability, a logarithmic scale was applied to the horizontal axis. During the training process, the bias correction for CR was implemented when recording return values to enable direct comparison with GR’s training progression. For DR, due to the inability to align return values with GR and CR under a unified metric through simple transformations, dual y-axes were employed for data logging. Consequently, GR and CR share the right y-axis, facilitating comparisons of absolute return values and training trend. The training curves of all 3 algorithms are plotted in a single figure to primarily compare convergence speed and stability. It is noteworthy that when presenting the final graphical results, we aligned the left and right axes based on the quality of the converged policies. Specifically, if a converged curve appears higher, we ensured that statistically significant return test results would be larger within the displayed range. Due to the nonlinear numerical mapping between differential rewards and the other two reward methods, relative relationships may exhibit inconsistencies across the entire vertical coordinate range. Quantitative performance evaluations of the algorithms are conducted during the model testing phase.
@@
It can be observed in Fig. 3 that in MADQN, the training curves for all three reward functions exhibit significant fluctuations, indicating poor stability. Despite tuning the algorithm to its optimal state, the performance of the three reward functions varies across different penetration rates. This instability is likely due to MADQN being based on independent Q-learning, which struggles to adapt to multi-vehicle cooperative decision-making tasks. We will discuss this phenomenon in detail in the later sections. Additionally, value-iteration algorithm like DQN have more reward variance than another rewards, which may also lead to the fluctuations. The high variance of value iteration methods also manifests in the QMIX implementation. In MAPPO, although the algorithm achieves stable convergence, the performance of the centralized reward function exhibits considerable deviation in the stabilized phase under the same settings. Specifically, for penetration rates of 0.25, 0.5, and 1.0, the orange curve shows substantial deviation after 20,000 episodes.
@@
It is evident that the convergence speed of the differentiated reward function is significantly faster than that of the centralized and general reward functions, especially in experiments with MAPPO and QMIX. Additionally, the return values of the centralized reward function are noticeably higher than those of the general reward function, particularly in the later training stages (¿ 20,000 episodes). This finding is consistent with the conclusions reported in [19].
Table III shows the performance metrics of the 3 reward functions under the QMIX algorithm at different penetration rates. For each algorithm under every simulation configuration, we conducted three independent trials, with each trial comprising 1000 episodes and utilizing the optimal model obtained in the training phase.
@@
It can be observed that in most cases the differentiated reward function achieves the best performance in terms of traffic efficiency, safety, action rationality, and task completion rate.
The DR maintains relatively high speeds (11.46-16.79 m/s) and safe spacing (32.30-53.56 m) across Pene. Rate of 25%-100%, indicating its ability to balance efficiency and safety. At 50% Pene. Rate, CR achieves a higher average speed (14.18 m/s) compared to DR (12.06 m/s), but demonstrates a significantly smaller Min. Gap of 5.55 m versus DR’s 49.04 m. This suggests CR prioritizes speed enhancement while tolerating reduced safety margins, whereas DR effectively suppresses risky overtaking behaviors through positional reward. Under 75% and 100% Pene. Rate, GR exhibits notably lower average speeds compared to normative levels, reflecting deficiencies in its reward design that induce vehicular stagnation phenomena, potentially due to excessive penalization of lane changes or conflicting intention resolution mechanisms.
Regarding the LC. Freq., DR exhibits the lowest values except 100% Pene. Rate, demonstrating that the positional reward in Equation (23) effectively guides vehicles to stabilize in target lanes promptly, thereby reducing unnecessary lane changes. At 25%, 50%, and 75% penetration rates, GR and CR show significantly higher lane-changing frequencies than DR. This is hypothesized to result from smaller penalty weights on lane changes, where reward for lane-changing are frequently overshadowed by value function estimation errors, leading to inefficient maneuvers. Under the high penetration rate (100%) scenario, both CR and DR display notable increases in LC. Freq., which through analysis of test results is attributed to elevated CAV density, which induces reciprocal displacement among vehicles during interactive processes.
Tables IV and V present the test results of MADQN and MAPPO respectively. The MADQN test data (Table 4) shows that, as indicated by the training curves, MADQN’s overall performance is limited, with generally low Succ. Rate across all reward functions. This further confirms the inherent limitations of independent Q-learning-based MADQN in effectively handling multi-agent collaborative tasks. Nevertheless, the DR method demonstrates significant advantages in several key metrics. The most notable improvement lies in safety: under all penetration rates, the Min. Gap with DR policies is substantially greater than those of GR and CR. For instance, at a 25% penetration rate, DR achieves a minimum distance of 17.00m, compared to the dangerously small 1.85m and 4.05m for GR and CR respectively. This indicates that even when the base algorithm is unstable, DR’s reward mechanism can effectively suppress risky behaviors. Additionally, DR consistently outperforms GR and CR in success rates, suggesting that DR can guide policy optimization toward more effective directions, despite the ultimate performance being constrained by the algorithm itself.
In Table 5, compared with MADQN, MAPPO demonstrates significantly improved overall performance, but the performance differences between reward functions become more pronounced. The DR function comprehensively and substantially outperforms GR and CR across all four penetration rates and all four performance metrics. Particularly in terms of average speed and success rate, DR’s advantages are most evident. For example, at 100% penetration rate, DR achieves an average speed of 18.47 m/s and a success rate of 66.93%, while CR and GR exhibit poor performance.
An intriguing observation concerns the lane change frequency. Under the MAPPO algorithm, the lane change frequency with the DR strategy is extremely low, significantly lower than that of GR and CR (except at 100% penetration rate). This strongly demonstrates that the positional reward in DR effectively guides vehicles to stabilize quickly after reaching the target lane, reducing unnecessary or high-risk lane changes caused by inaccurate value function estimation. This result also explains why DR can maintain high traffic efficiency while achieving excellent minimum gap and success rates.
In summary, for different types of reinforcement learning algorithms, the reward differential method can ensure fundamental safety and improve task success rates. Under the QMIX algorithm, which exhibits superior overall performance, DR achieves success rates above 88% across penetration rates of 25%-100%, significantly outperforming CR and GR. The positional reward $r_{p}^{i}$ in DR is dynamically adjusted through gradient $\nabla f_{p}^{i}$, enabling vehicles to obtain higher rewards near target lanes. The algorithm involving the dot product of velocity and positional gradients, as shown in Equation (22), further guides vehicles to mitigate failures caused by trajectory deviations.
@@ VI Conclusion and future work
This paper proposes a differentiated reward method for reinforcement learning-based multi-vehicle cooperative decision-making algorithms. By incorporating gradient information of state transitions into reward function design, this approach addresses the limitations of conventional reward mechanisms in distinguishing action values within steady-state traffic flows. Experimental results demonstrate that the proposed method significantly enhances the convergence speed and training stability of multi-agent reinforcement learning algorithms. Strategies derived with the proposed reward function also outperforms those from centralized reward functions and general reward shaping methods across core metrics including traffic throughput, safety metrics, and action rationality. Furthermore, the differentiated reward method maintains stable performance across varying autonomous vehicle penetration rates, exhibiting robust scalability.
However, the differentiated reward method has limitations. Firstly, in value-iteration-type reinforcement learning algorithms, the gradient of the reward function amplifies action-value discrepancies, potentially inducing higher reward variance as evidenced in Fig. 3 (such issues also exist in centering reward functions but manifest more severely in the reward-differential framework). This phenomenon may degrade overall algorithm performance. Additionally, our current implementation uses a discrete action space, which simplifies the gradient computation. Extending this to continuous control scenarios would require calculating higher-order vector derivatives of the reward function, posing a significant computational challenge.
Future work will focus on refining the theoretical foundations of differentiated reward function design to accommodate more reinforcement learning applications. Generalizability to More Complex Scenarios is also important. Our experiments were conducted on a multi-lane highway, which raises questions about the method’s applicability to more complex traffic environments. We plan to extend the DR framework to challenging scenarios such as unsignalized intersections, roundabouts, and merging zones. This will involve adapting the potential field’s concept from a static ”target lane” to a dynamic series of waypoints derived from trajectory planning, thus accommodating more fluid and conflicting vehicle objectives. Additionally, the current work demonstrates adaptability to varying CAV penetration rates but does not cover other critical environmental variables. A key priority will be to evaluate the method’s performance under a wider spectrum of traffic conditions, from free-flow to congested traffic states, and under different traffic flow distributions. Furthermore, we will assess the algorithm’s robustness against adversarial disturbances, such as sudden pedestrian intrusions or simulated sensor noise, to better evaluate its readiness for real-world deployment.